A point moves along a straight path. The function
f(t)=log3(t)
determines the distance (in meters) the point has traveled in terms of the number of seconds
t
since the point started moving.
How far has the point traveled 22 seconds after it started moving?
meters
If the point has traveled 2 meters, how many seconds have elapsed since it started moving?
seconds
Write a function
f^-1
that determines the number of seconds that have elapsed since the particle started moving in terms of the distance (in meters) the particle has traveled,
d.f^−1(d)=


Answer by Boreal(15235)

when t=22, the point has moved log 66 m or 1.819 m
2= log 3 t
raise to 10th power
100=3t
t=100/3 seconds
The inverse function is t=log (3y
10^t=3y
y=(1/3)(10^t)
NOTE: I am assuming this is log 3t not log(3) t